OpenStudy (astrophysics):

@ganeshie8 @dan815 @Kainui

OpenStudy (astrophysics):

I don't understand how to use coordinates if it's asking for linear dependence, because I need a basis and that would require it to be linearly independent. Huh 

OpenStudy (astrophysics):

\[c_1p(x)+c_2q(x)+c_3r(x) = 0 \]

OpenStudy (dan815):

okay so u got 1,-1-1 as 1 solution

OpenStudy (dan815):

so uve shown lin dependance

OpenStudy (astrophysics):

Not yet I was doing that right now, but I can doing what I am above, one second, \[c_1(2-x+x^2)+c_2(1+3x+3x^2)+c_3(1-4x-2x^2)=0\]

OpenStudy (astrophysics):

\[2c_1-xc_1+c_1x^2+c_2+3xc_2+3x^2c_2+c_3-4xc_3-2x^2c_3=0\] \[(2c_1+c_2+c_3)+(-c_1+3c_2-4c_3)x+(c_1+3c_2-2c_3)x^2=0\] After equating the sides we get \(2c_1+c_2+c_3=0\) \(-c_1+3c_2-4c_3=0\) \(c_1+3c_2-2c_3=0\)

OpenStudy (dan815):

mhm

OpenStudy (astrophysics):

Did I make a mistake

OpenStudy (dan815):

no its fine

OpenStudy (dan815):

ull get a dependance saying c2=c3 c1=c2+c3

OpenStudy (astrophysics):

Right

OpenStudy (dan815):

okey

OpenStudy (dan815):

go onnnn

OpenStudy (astrophysics):

I thought i had to use coordinates on every part of the question xD so i was like wut

OpenStudy (dan815):

just ignore the coordinate word for now

OpenStudy (dan815):

solve it like u normally wud

OpenStudy (astrophysics):

Ok, so for a dependence relation I'd plug in the values for c1 c2 c3, so for now I'll just use what you said, of course I'll do all the technical work after, so c1=1, c2=-1, c3=-1 \[1 \times p(x)-1 \times q(x)-1 \times r(x) = 0\]

OpenStudy (dan815):

ya

OpenStudy (dan815):

because u can see right like if u add q(x) + r(x) u get p(x)

OpenStudy (astrophysics):

Yes

OpenStudy (dan815):

p=q+r 1p=1q+1r 1p-1q-1r=0 so 1,-1,-1 is a solution

OpenStudy (astrophysics):

Omg >_>

OpenStudy (dan815):

similarily k*<1,-1,-1> , k E R is a solution

OpenStudy (dan815):

this solves a,b

OpenStudy (dan815):

now what is the definition of span

OpenStudy (astrophysics):

All the linear combinations possible in the set

OpenStudy (astrophysics):

the set of vectors in V

OpenStudy (dan815):

okay so just take q and p

OpenStudy (dan815):

that is one basis

OpenStudy (astrophysics):

So I don't have to show they span S right as that's a given, I must show it's linearly indep however?

OpenStudy (astrophysics):

Because a basis must span and be lin. indp.

OpenStudy (dan815):

right

OpenStudy (astrophysics):

Ok

OpenStudy (dan815):

u can take any 2 of {q,p,r}

OpenStudy (astrophysics):

So I can show its lin indep, is there a technique like the one before haha or do I have to do it the long way as I was before

OpenStudy (astrophysics):

I'd use p and q to be safe

OpenStudy (dan815):

okay well since its just 2 vectors in this case

OpenStudy (dan815):

since one is not a scalar multiple u can already say its lin independant

OpenStudy (astrophysics):

Eh I'd rather show it though, like \[c_1p(x)+c_2q(x)=0\] would that work?

OpenStudy (dan815):

ya u can do that too

OpenStudy (astrophysics):

Ok cool, that will be easy to since it's only 2 vectors \[c_1(2-x+x^2)+c_2(1+3x+3x^2)=0\] \[2c_1-xc_1+x^2c_1+c_2+3xc_2+3x^2c_2=0\] \[(2c_1+c_2)+(-c_1+3c_2)x+(c_1+3c_2)=0\] \(2c_1+c_2=0\) \(-c_1+3c_2=0\) \(c_1+3c_2=0\) Yeah it'll be linearly independent

OpenStudy (astrophysics):

So is our basis just \[S = \left\{ p(x),q(x) \right\}\] as well

OpenStudy (dan815):

ya

OpenStudy (dan815):

u have an finite number of basis thats just one

OpenStudy (dan815):

infinite*

OpenStudy (astrophysics):

Ok I see, these terms are annoying at times lol

OpenStudy (astrophysics):

So then c is done

OpenStudy (dan815):

ya now try solving c1p+c2q=f

OpenStudy (astrophysics):

For \[f(x)=x+3x^2\] now I must check if it spans and is lin indep

OpenStudy (astrophysics):

in S

OpenStudy (dan815):

f(x) doesnt span anything by itself its just 1 "vector"

OpenStudy (dan815):

nono u gotta solve

OpenStudy (dan815):

see if ur basis can hit that vector

OpenStudy (dan815):

if a linear combination exists then f(x) is in S

OpenStudy (dan815):

i think u cant do it

OpenStudy (astrophysics):

Right, so \[f(x) = c_1p(x)+c_2q(x)\]

OpenStudy (astrophysics):

Let me try

OpenStudy (dan815):

okay

OpenStudy (dan815):

i found a trick to doing these without writing equations

OpenStudy (astrophysics):

\[x+3x^2=c_1(2-x+x^2)+c_2(3x+3x^2)\] Really, because I'm tired of equating sides lol

OpenStudy (astrophysics):

What are you talking about

OpenStudy (astrophysics):

Did you discover a right hand rule for this

OpenStudy (astrophysics):

xD

OpenStudy (astrophysics):

Ok

OpenStudy (astrophysics):

I'm going to write it on draw, cause equation takes so long

OpenStudy (dan815):

brb getting food

OpenStudy (astrophysics):